OpenStudy (anonymous):

let z=z(x,y) be given imlicitely by sin(x+y)+cos(x+y)=pi/4 find Zxy..

myininaya (myininaya):

I don't get what Zxy means

myininaya (myininaya):

Can you tell me what it means please?

OpenStudy (anonymous):

it is derivative of Z wrt. x first and second derivative of y

myininaya (myininaya):

oh I think I understand sorry

myininaya (myininaya):

So we need Zx first!

myininaya (myininaya):

\[(x+y)_xcos(x+y)+(x+y)_x(-\sin(x+y))=0\] \[(1+0)\cos(x+y)+(1+0)(-\sin(x+y))=0\] \[1\cos(x+y)+1(-\sin(x+y))=0\] \[\cos(x+y)-\sin(x+y)=0\] Since I was taking the partial derivative w.r.t. x I treated y as a constant Do you see that?

myininaya (myininaya):

So that is Zx

myininaya (myininaya):

We need to do (Zx)y now

myininaya (myininaya):

So we will look at Zx and we will take partial derivative of it with respect to y that means we will treat x like it is a constant

myininaya (myininaya):

Do you want to try?

OpenStudy (anonymous):

-cos(x+y)-sin(x+y)=0 right ?

myininaya (myininaya):

\[(x+y)_y(-\sin(x+y))-(x+y)_ycos(x+y)=0\] \[(0+1)(-\sin(x+y)-(0+1)\cos(x+y)=0\] \[-\sin(x+y)-\cos(x+y)=0\] yes that is right that is Zxy